# Navigating the Future of Federated Recommendation Systems with Foundation Models

### v2

@@ Abstract
In recent years, the integration of federated learning (FL) and recommendation systems (RS), known as Federated Recommendation Systems (FRS), has attracted attention for preserving user privacy by keeping private data on client devices. However, FRS faces inherent limitations such as data heterogeneity and scarcity, due to the privacy requirements of FL and the typical data sparsity issues of RSs. Models like ChatGPT are empowered by the concept of transfer learning and self-supervised learning, so they can be easily applied to the downstream tasks after fine-tuning or prompting. These models, so-called Foundation Models (FM), fouce on understanding the human’s intent and perform following their designed roles in the specific tasks, which are widely recognized for producing high-quality content in the image and language domains. Thus, the achievements of FMs inspire the design of FRS and suggest a promising research direction: integrating foundation models to address the above limitations. In this study, we conduct a comprehensive review of FRSs with FMs. Specifically, we: 1) summarise the common approaches of current FRSs and FMs; 2) review the challenges posed by FRSs and FMs; 3) discuss potential future research directions; and 4) introduce some common benchmarks and evaluation metrics in the FRS field. We hope that this position paper provides the necessary background and guidance to explore this interesting and emerging topic.
@@
Federated Recommendation System, Foundation Model, Privacy Preserving, Security,
@@ I Introduction
In the digital age, the exponential growth of information has created a need for systems that navigate, filter, and personalize data for individual users. Recommendation Systems (RS) nowadays become crucial tools for filtering online information and helping users discover products, content, and services that align with their preferences [1]. However, the systems traditionally rely heavily on centralized data collection and processing, posing significant privacy risks and operational bottlenecks. The importance of user privacy has never been greater, particularly with stringent data protection regulations such as the General Data Protection Regulation (GDPR) [2] in Europe, which emphasises the need to store user data on theirs local devices, instead of uploading it to central servers. As a novel approach to address these privacy concerns, Google introduced Federated Learning (FL) [3] as a framework designed to train models across decentralized devices, while keeping data localized. This paradigm shift in data processing leverages the computational capabilities of individual devices for local data analysis. Specifically, FL alternates between local model training at the client and global parameter aggregation from these models on a central server. The integration of FL with RS becomes essential for safeguarding user privacy in recommendation services, which has given rise to the burgeoning field of Federated Recommendation Systems (FRS) [4]. In this field, typically each client consists of a single user’s device. Therefore, unless specified otherwise in this paper, the terms ’user,’ ’client,’ and ’device’ all refer to an individual user. FRS recently have shown promising results in many areas, such as service providing [5, 6], daily scheduling [7], driving planing [8, 9] and more, significantly impacting different facets of daily life. Similar to FL, FRS is required to keep user data local to protect user privacy. However, in most cases, each client typically contains only the data of one user’s accessed item, which is extremely small compared to the total number of item set, creating a serious data scarcity problem. In addition, different users have different behaviours and preferences, which can lead to data heterogeneity. The presence of both issues can lead to sub-optimal models and reduced effectiveness.
@@
More recently, the emergence of a novel paradigm for the construction of artificial intelligence (AI) systems has garnered considerable interest in the wake of the remarkable success of ChatGPT [10] and diffusion models [11, 12, 13] in the tasks of language understanding [14, 15, 16, 17] and image generation [12, 13], which we refer to as Foundation Models (FM) [18]. As shown in Fig. 1, this paradigm is built by using self-supervised optimization of training goals to determine how to update model parameters based on model predictions on the large amount of unlabelled training data. This process is referred to as pre-training. Then, FM can utilize a technique, called Adaptation, to turn the general FM into a task-specific one. This process is termed as Fine-tuning or Prompting. Language models, e.g. BERT [19]and RoBERTa [20], are usually trained using the next token prediction goal, which refers to the extent to which the model is able to predict the next token in the sequence. One of the most successful examples of language models is ChatGPT, which is based on GPT-3.5. By training on large amounts of text data, it aligns the capabilities of large language models (LLM) with human intent [18]. Vision models like ViT [21] are typically trained using either contrast learning or diffusion training targets. For contrastive learning [22], images are randomly augmented before evaluating the similarity of the model representations. For diffusion models [11, 12, 13], noise is added to the images and the model is gradually de-noised by the target. There are also multi-modal training targets, some of which separate images and text during training, while others consider both. The foundation model, once trained, is used as a plug-in in combination with adaptations to achieve results for a wide variety of downstream tasks. CLIP [23] and DALL·E [24] are two multi-modal AI models developed by OpenAI. CLIP focuses on understanding images through natural language, while DALL·E focuses on generating images based on text descriptions. While both models are trained on large datasets of images and text, CLIP is primarily used for image retrieval and classification, while DALL·E excels at generating new images that match text descriptions [25].
@@
As FM is frequently trained using a substantial quantity of data from multiple sources, they are capable of incorporating a considerable amount of additional knowledge when performing a specific downstream task. This feature enables FM to provide diverse training data for task-specific models in downstream task scenarios, thus effectively alleviating the problem of data scarcity. Given these considerations, the application of FM to FRS is not only a viable approach but also holds significant research potential. Such integration will drive innovation in FRS. Although the application of FM to FRS has the potential to be highly beneficial, it must be acknowledged that this field is still in its nascent stages, with an insufficient understanding of the challenges, viable methods, and directions for development. This paper aims to bridge this knowledge gap through an in-depth analysis of the integration of FM and FRS. The article provides a comprehensive examination of the motivations and challenges associated with integrating these two paradigms, with a particular focus on several representative technologies. Additionally, it outlines future development trends and their applications. By elucidating the intersection of FRS and FM, this study aims to promote further exploration and innovation within this emerging field, thereby facilitating its rapid advancement.
II Related Work and Contribution
@@
In the field of FRSs, various studies have converged on methodologies, privacy preservation, and challenges, albeit with distinct focal points. Yang et al. [26] discuss the practical implementation and evaluation of FRS, with a focus on system architecture and algorithmic efficiency. Alamgir et al. [27] provides a comprehensive overview of FRS, highlighting techniques, prevailing challenges, and future directions. Javeed et al. [28] focus on the security and privacy concerns in personalized recommendation systems and propose targeted solutions to these challenges. Finally, Sun et al. [29] conduct a survey that compares existing federated recommendation research, highlighting the strengths and limitations of various approaches. All studies emphasize the critical importance of privacy protection and the challenges posed by data heterogeneity and model aggregation. This body of work enhances our understanding of FRS and provides a foundation for future research to address their inherent challenges. Although a number of study focusing on combining FM with FL [30, 31, 32, 33, 34, 35, 36] or RS [37, 38] already exist, our paper is the first work to explore the integration of FM into FRS.
Contribution. This survey aims to provide a clear theoretical framework for applying FM to FRS and to elaborate on their principles and methods of application. By analyzing existing technologies, this paper strives to drive innovation in integrating the pre-training capabilities of FM into FRS. The article discusses in detail the technical challenges and practical issues faced when integrating FM, such as privacy protection, data heterogeneity, communication efficiency, and model generalization capabilities. It also identifies current research gaps and future directions, aiming to guide subsequent academic research and technological development. Additionally, through application studies, this paper demonstrates how to apply the integration of FM with FRS in practical scenarios, thereby deepening the integration of theory and practice.
@@ III-A Federated Recommendation System
Figure 2: A typical framework for FRS.
@@
Figure 3: Typical data categories of FRS.
Definition 1 (Federated Recommendation System)
FRS facilitate collaborative learning by pooling analysis and learning capabilities across multiple clients (e.g. users or locations), without the need for direct exchange of raw user data. Fig. 2 illustrates that this type of system typically involves three processes: 1) client model update, 2) communication, and 3) global aggregation. The client model update allows users to train their models locally on their devices using their own data and then upload the updated intermediate parameters, e.g. the model gradients or parameters can be shared, to the server. The server then performs global aggregation on the parameters sent by all participants, integrating the unique information from each of them. The aggregated parameters are then distributed to the next round of participants, initiating a new training round. The process of uploading and distributing is collectively referred to as communication. These approaches effectively preserves privacy and security.
III-A1 Taxonomy of FRS
Figure 4: The diagram illustrates the taxonomy of FRS based on two main criteria: Data Distribution and Communication Architecture.
As illustrated in Fig. 4, FRS can be categorized based on different criteria [26, 29]. According to used data distribution, FRS can be divided into Horizontal FRS, Vertical FRS, and Transfer FRS. Each type addresses different data collaboration and learning scenarios. Specifically:
  • Horizontal FRS deals with situations where the feature spaces are similar across different entities, but the sample spaces differ, as shown in Fig. 3a. For instance, two customers who shop at the same store may purchase some of the same products. The horizontal method aggregates model updates from various sources to train recommendation models without sharing raw user data. This approach enhances the accuracy and efficiency of recommendation systems while safeguarding user privacy. Horizontal FRS is currently the most common type, with many studies [39, 40, 41] based on the assumption of this data distribution.
  • Vertical FRS is applied when the feature spaces differ, but the sample spaces are similar, as shown in Fig. 3b. This scenario is common when different parties possess different types of data about the same users. An example would be a bank and an online retailer holding distinct perspectives on the same customers, i.e., the bank has credit history while the retailer has shopping history. One challenge is that how to align the identities of customers across different parties. Vertical FRS [42, 43, 44, 45] trains recommendation models by securely integrating different data features between parties, utilizing richer user information for more accurate personalized recommendations and ensuring data privacy and security.
  • Transfer FRS utilizes principles of transfer learning [46] to transfer knowledge from one domain to another, as shown in Fig. 3c. This type is suited for scenarios with significant differences in both feature and sample spaces across parties. Transfer learning allows the system to leverage data and knowledge from the source domain to enhance recommendation performance in the target domain, even when the target domain lacks sufficient data for independent model training. Transfer FRS [47] is particularly beneficial for emerging markets or user groups with limited data, drawing insights from related domains.
@@ III-A1 Taxonomy of FRS
Figure 5: Typical architecture categories of FRS.
As shown in Fig. 5, FRS can be categorized into centralized, semi-decentralized and decentralized based on their communication architecture. Each type addresses privacy and scalability in different ways:
  • Centralized FRS operates within a FL framework where a central server orchestrates the learning process as shown in Fig. 5a. Clients locally compute updates based on their data and send these updates to the server. The server aggregates these updates to prompt essential information sharing, which is then distributed back to the users. It effectively addresses privacy concerns by allowing the model to learn from decentralized data sources without centralizing the data itself. While this approach improves privacy by not requiring the sharing of raw data, it still relies on a central authority to manage the model. This architecture is typically used to overcome data silo issues [48] and improve the performance of RS without compromising user privacy and data security. Due to its simplicity and intuitive nature, this architecture has become the dominant framework within the field of FRS. There is a great deal of work [49, 39, 40, 41, 47] based on it currently.
  • Semi-decentralized FRS shown in Fig. 5b introduces an intermediate layer between the central server and the users, such as edge servers or devices that can perform additional computational or storage tasks. This setup aims to reduce the communication overhead and latency associated with sending updates to a central server, especially in large-scale applications. A specific example is the SemiDFEGL framework [50], which improves scalability and reduces communication costs by introducing new device-to-device collaborations. It augments local subgraphs with predicted interacted item nodes to exploit high-order collaborative information between users and items in a privacy-preserving manner, which is particularly useful for recommendations based on collaborative filtering and graph neural networks.
  • Decentralized FRS, which employs peer-to-peer communication architecture [51], distributes the learning process completely to all participating devices, without the need for a central server for model aggregation, as shown in Fig. 5c. This approach maximizes privacy and data ownership, but presents challenges in coordinating model updates and ensuring model convergence. Zheng et al. [52] proposed a FRS model called DGREC by adopting a decentralized graph neural network, which constructs a local intra-item hypergraph and a global inter-user graph for each user, allowing users to freely choose whether to disclose their interactions. Li et al. [53] introduced DFedRec, which uses a privacy-aware client-level structured graph to share model parameters only with relevant neighboring users, thereby reducing communication costs and protecting user privacy.
@@ III-B Foundation Model
The rapid increase in the performance of computer hardwares, e.g., GPUs, the increasing maturity of transformer architectures, and the public availability of large amounts of training data have been three key factors in the emergence of FM [18]. According to the work of Stanford HAI [18], we have the following definition of FM:
@@
In recent years, the scale and scope of FM have greatly expanded our imagination of potential applications. Such models typically have billions or even trillions of parameters, allowing them to learn more complex patterns and knowledge. They can be adapted to new tasks through fine-tuning or prompting [54, 55]. For instance, the GPT-3 model [56], comprising 175 billion parameters, is capable of adapting and perform various tasks with the aid of natural language prompts, despite the fact that a significant proportion of these tasks have not been explicitly trained. FM is distinguished by two key characteristics [18]: 1) Emergence, which refers to the implicit induction of system behavior from examples, as opposed to explicit design; and 2) Homogenization, which indicates that the method of constructing machine learning systems tends to be unified across a wide range of applications. Although FM is based on self-supervised learning and transfer learning, their scale brings new emergent abilities, and their effectiveness in many tasks has motivated homogenization, e.g., many FMs have evolved based on transformer architectures [19, 20, 21, 56]. While homogenization provides strong leverage, it also requires caution, as defects in foundation models can be inherited by all downstream adapted models. Despite the broad deployment prospects of foundation models, due to their emergent properties, our understanding of how they work, when they fail, and what they can actually do is still quite limited [18], but there has been quite a bit of work [57, 58, 59] that attempts to explain this phenomenon. Moreover, OpenAI proposes a Scaling Law for LLM [60], suggesting that the performance of a model is governed by a power-law relationship involving computational power, model parameters, and data size. This law also applies to downstream tasks [61]. Therefore, understanding and leveraging this power-law relationship when using FM can aid in more efficiently allocating resources and optimizing model performance.
Figure 6: The diagram illustrates the taxonomy of FM based on two main criteria: Data Type and Functionality.

### v3

@@ Abstract
In recent years, the integration of federated learning (FL) and recommendation systems (RS), known as Federated Recommendation Systems (FRS), has attracted attention for preserving user privacy by keeping private data on client devices. However, FRSs face inherent limitations such as data heterogeneity and scarcity, due to the privacy requirements of FL and the typical data sparsity issues of RSs. Models like ChatGPT are empowered by the concept of transfer learning and self-supervised learning, so they can be easily applied to the downstream tasks after fine-tuning or prompting. These models, so-called Foundation Models (FM), focus on understanding the human’s intent and perform following their designed roles in the specific tasks, which are widely recognized for producing high-quality content in the image and language domains. Thus, the achievements of FMs inspire the design of FRS and suggest a promising research direction: integrating foundation models to address the above limitations. In this study, we conduct a comprehensive review of FRSs with FMs. Specifically, we: 1) summarise the common approaches of current FRSs and FMs; 2) review the challenges posed by FRSs and FMs; 3) discuss potential future research directions; and 4) introduce some common benchmarks and evaluation metrics in the FRS field. We hope that this position paper provides the necessary background and guidance to explore this interesting and emerging topic.
@@
Federated Recommendation System, Foundation Model, Survey, Privacy Preserving, Security
@@ I Introduction
In the digital age, the exponential growth of information has created a need for systems that navigate, filter, and personalize data for individual users. Recommendation Systems (RS) have become crucial tools for filtering online information and helping users discover products, content, and services that align with their preferences [1]. However, these systems traditionally rely heavily on centralized data collection and processing, posing significant privacy risks and operational bottlenecks. User privacy has become increasingly important, especially with stringent data protection regulations such as the General Data Protection Regulation (GDPR) [2], emphasizing storing user data locally rather than uploading it to servers.
To address these privacy concerns, Google introduced Federated Learning (FL) [3], designed to train models across distributed devices while keeping data localized. This paradigm shift in data processing leverages the computational capabilities of individual devices for local data analysis. Specifically, FL alternates between training models locally on each client and aggregating parameters globally on a central server. The integration of FL with RS is essential for safeguarding user privacy in recommendation services, giving rise to the burgeoning field of Federated Recommendation Systems (FRSs) [4].
FRSs have shown promising results in various areas, such as service provision [5, 6], daily scheduling [7], and driving planning [8, 9], impacting multiple facets of daily life. However, in most cases, each client typically consists of a single user’s device, containing only the data of one user’s accessed items, which is extremely small compared to the total item set, creating a serious data scarcity problem. In addition, different users have different behaviors and preferences, leading to significant data heterogeneity. The presence of both issues can result in sub-optimal models and reduced effectiveness. Addressing these challenges is crucial for the success of FRSs, particularly as they become more widely adopted in privacy-sensitive domains. Innovative strategies to enhance data sharing and model personalization while maintaining privacy are needed to overcome data scarcity and heterogeneity.
@@
More recently, a novel paradigm for constructing artificial intelligence (AI) systems, known as Foundation Models (FMs) [10] (or Pretrained Large Models [11]), has garnered considerable interest following the remarkable success of models like ChatGPT [12] and various diffusion models [13, 14, 15]. As illustrated in Fig. 1, this paradigm involves using self-supervised learning objectives to pre-train models on large amounts of unlabeled data. This process, referred to as Pre-Training, enables the models to learn general-purpose representations. Once pre-trained, FMs can be adapted to specific tasks through techniques such as Fine-Tuning [16] or Prompting [17], turning the general model into a task-specific one. This adaptability allows FMs to achieve state-of-the-art performance across a wide range of applications. Examples of FMs include language models like BERT [18] and vision models like ViT [19], which have significantly advanced natural language understanding and computer vision, respectively. Multi-modal models like CLIP [20] and DALL·E [21] extend these capabilities to handle both text and images, enabling tasks such as image-text retrieval and text-to-image generation.
The widespread adoption and success of FMs demonstrate their profound impact on AI research and applications. FMs are typically trained on substantial quantities of data from diverse sources, enabling them to incorporate extensive knowledge when tackling specific downstream tasks [10]. This characteristic offers several significant benefits for FRSs [22]: Firstly, FMs can alleviate the data sparsity issue by providing rich pre-trained representations, thus enhancing recommendation accuracy even with limited user interactions. Secondly, they effectively mitigate the cold-start problem for new users and items by leveraging learned general patterns. Thirdly, FMs facilitate efficient transfer learning, allowing models to adapt quickly to new domains or tasks with minimal additional training. Moreover, the use of FMs can help balance privacy and performance; their pre-trained capabilities reduce the need for extensive data sharing and multiple rounds of communication in federated settings, thereby lowering the computational and communication costs associated with federated training. Given these considerable advantages, integrating FMs into FRSs is not just a viable approach but a necessary progression to overcome existing limitations and drive innovation in the field.
@@
Although integrating FMs into FRSs holds immense promise, it must be acknowledged that this field is still in its nascent stages, with an insufficient understanding of the challenges, viable methods, and directions for development. This paper aims to bridge this knowledge gap through an in-depth analysis of the integration of FMs and FRSs. We provide a comprehensive examination of the motivations and challenges associated with uniting these two paradigms, drawing on insights from recent studies, which highlight the benefits of introducing FMs in federated recommendation scenarios. Additionally, we outline future development trends and potential applications. By elucidating the intersection of FRSs and FMs, this paper aims to promote further exploration and innovation within this emerging field, thereby facilitating its rapid advancement. The paper also emphasizes the critical role of pre-trained representations in addressing challenges like data sparsity and heterogeneity, which are inherent in FRSs.
Paper Organization. The reminder is structured to provide a thorough study of the current research landscape, technical challenges, application cases, and future directions of FRSs. The organization is as follows: Sec. II outlines the existing surveys within the field of FRSs, and distinctly highlights our novel contributions, setting our paper apart from the current literature. Sec. III introduce foundational knowledge in the field of FRSs and FM, including definitions, basic architectures, and current scenario taxonomy. Sec. IV introduces the potential applications and functions of FM according to the three classical phases of FRSs. Sec. VII summarizes the datasets and evaluation methods commonly used in FRSs with FM, explaining how to quantify the recommendation performances. Sec. V delves into the main technical challenges faced by FRSs with FM, including data heterogeneity, model aggregation difficulties, and privacy concerns, as well as current methods and their limitations in addressing these challenges. Sec. VI explores future research directions in FRSs, including unresolved issues and emerging research areas, providing guidance and inspiration for researchers in the field. Sec. VIII concludes our paper with a summary of key findings and perspectives, offering insights into the trends and potential future of FRSs.
II Related Surveys and Contribution
@@
In the field of FRSs, several surveys have been conducted to synthesize methodologies, privacy preservation techniques, and inherent challenges, each emphasizing different aspects. Yang et al. [23] discuss the practical implementation and evaluation of FRSs, focusing on system architecture and algorithmic efficiency. Alamgir et al. [24] provide a comprehensive overview of FRSs, highlighting techniques, prevailing challenges, and future directions. Javeed et al. [25] focus on security and privacy concerns in personalized RSs, proposing targeted solutions to these challenges. Finally, Sun et al. [26] conduct a survey comparing existing federated recommendation research, highlighting the strengths and limitations of various approaches. Collectively, these surveys emphasize the critical importance of privacy protection and address challenges posed by data heterogeneity and model aggregation in FRSs. This work enhances our understanding of FRSs and provides a foundation for future research to tackle challenges.
While several studies have explored the combination of FMs with FL [27, 28, 29, 30, 31, 32, 33] or RSs [34, 35], to the best of our knowledge, no existing work has systematically analyzed or discussed the integration of FMs into FRSs. Our paper aims to fill this gap by providing a comprehensive examination of this intersection.
Contributions. Our main contributions are as follows:
  • We establish a comprehensive framework for applying FMs to FRSs, elaborating on the underlying principles and methodologies for effective integration.
  • By thoroughly analyzing existing technologies, we inspire innovation in leveraging the pre-training capabilities of FMs within FRSs, highlighting potential advancements.
  • We delve into the challenges and practical issues of integrating FMs into FRSs, including privacy protection, data heterogeneity, communication efficiency, and model generalization, providing insights and potential solutions.
  • We identify current research gaps and outline future directions, aiming to guide subsequent academic research and technological development in this emerging area.
  • Through application studies, we illustrate how the integration of FMs with FRSs can be applied in real-world scenarios, bridging the gap between theory and practice.
By elucidating the integration of FMs and FRSs, our paper seeks to promote further exploration and innovation, facilitating rapid advancement and broader adoption in the field.
Figure 2: The taxonomy of FRS and FM frameworks categorized by their respective core criteria in Sec. III.
@@ III Background and Preliminary
To provide a comprehensive understanding of the integration of FRSs and FMs, this section outlines their core principles, taxonomies, and foundational structures, as illustrated in Fig. 2, establishing a basis for analyzing their synergies.
@@ III-A Federated Recommendation System
Figure 3: A framework for FRSs, illustrating client-side local training with private data, server-side global aggregation, and update communication. Local models integrate user and item embeddings with prediction networks, while the server aggregates the sent local updates to enhance the global model, ensuring privacy and personalization.
Definition 1 (Federated Recommendation System)
FRSs enable collaborative learning across multiple clients without direct exchange of raw user data. Fig. 3 illustrates that FRS typically involves three key processes: 1) client model update, 2) communication, and 3) global aggregation. Users train models locally using their own data and share updated parameters (e.g., gradients) with a server. The server then performs global aggregation, integrating information from all participants and redistributing updated parameters. This communication process preserves privacy and security.
@@
Figure 4: Typical data distribution paradigms in FRS: (a) Horizontal FRS, characterized by similar feature spaces across users but distinct data samples; (b) Vertical FRS, involving overlapping user sets with disparate feature spaces; and (c) Transfer FRS, leveraging transfer learning to enable knowledge sharing across domains with limited or complementary data distributions.
As illustrated in Fig. 2, FRSs can be categorized based on data distribution and communication architecture [23, 26].
III-A1 Data Distribution Paradigms in FRS
There are three categories based the data distribution, each addresses specific data collaboration and learning scenarios:
  • Horizontal FRS deals with similar feature spaces but different sample spaces, as shown in Fig. 4a. It aggregates model updates from multiple clients to train recommendation models while preserving privacy [36, 37, 38].
  • Vertical FRS is applied when feature spaces differ but sample spaces are similar, as shown in Fig. 4b. It securely integrates different data types (e.g., a bank’s credit history) to enrich user information [39, 40, 41, 42].
  • Transfer FRS utilizes transfer learning [43] to transfer knowledge from one domain to another, benefiting scenarios with limited data for independent model training [22], shown in Fig. 4c.
Each type of FRS based on data distribution is designed to address specific collaboration scenarios, ensuring privacy and efficient learning. Horizontal FRS is suitable for situations where similar features are available across clients, Vertical FRS integrates different types of features for the same users, and Transfer FRS allows knowledge transfer across domains, making it particularly useful when data is scarce.
@@ III-A1 Data Distribution Paradigms in FRS
Figure 5: Architectural paradigms of FRS: (a) Centralized FRS, relying on a central server for model aggregation; (b) Semi-decentralized FRS, incorporating intermediate nodes to balance scalability and efficiency; and (c) Decentralized FRS, enabling peer-to-peer communication to enhance privacy and robustness.
III-A2 Communication Architectures in FRS
As shown in Fig. 5, Based on communication architecture, FRS can be categorized into centralized, semi-decentralized and decentralized based on their communication architecture. Each type addresses privacy and scalability in different ways:
  • Centralized FRS: A central server orchestrates the learning process in Fig. 5a, simplifying model aggregation to overcome data silo issues [44] while relying on a central authority. This approach is straightforward but depends heavily on the central server, creating a single point of failure and potential privacy risks [45, 36, 37, 38, 22].
  • Semi-decentralized FRS: As shown in Fig. 5b, It introduces an intermediate layer (e.g., edge servers) to reduce communication overhead and enhance scalability. This architecture balances the advantages of centralized and decentralized systems, improving system efficiency while still managing privacy concerns [46].
  • Decentralized FRS: It employs peer-to-peer communication [47], which maximizes privacy but necessitates complex coordination. As illustrated in Fig. 5c, each client communicates directly with others without relying on a central server. This enhances privacy and reduces dependence on a single entity; however, it comes at the expense of increased communication complexity and the requirement for robust consensus mechanisms [48, 49].
@@ III-B Foundation Model
The rapid increase in the performance of computer hardwares, e.g., GPUs, the increasing maturity of transformer architectures, and the public availability of large amounts of training data have been three key factors in the emergence of FMs [10]. According to the work of Stanford HAI [10], we have the following definition of FMs:
@@
FMs are pre-trained on extensive data, enabling adaptation to a wide range of downstream tasks through fine-tuning or prompting [50, 51]. These models, e.g. GPT-3 [52], are distinguished by two key characteristics [10]: 1) Emergence, refers to the model’s ability to develop new capabilities from training data without explicit programming; and 2) Homogenization, highlights the unified approach across diverse machine learning tasks, making FMs adaptable across multiple domains.
Although FMs are based on self-supervised learning and transfer learning, their scale brings new emergent abilities, and their effectiveness in many tasks has motivated homogenization, e.g., many FMs have evolved based on transformer architectures [18, 53, 19, 52]. While homogenization provides strong leverage, it also requires caution, as defects in foundation models can be inherited by all downstream adapted models. Despite the broad deployment prospects of foundation models, due to their emergent properties, our understanding of how they work, when they fail, and what they can actually do is still quite limited [10], but there has been quite a bit of work [54, 55, 56] that attempts to explain this phenomenon. Moreover, OpenAI proposes a Scaling Law for LLM [57], suggesting that the performance of a model is governed by a power-law relationship involving computational power, model parameters, and data size. This law also applies to downstream tasks [58]. Therefore, understanding and leveraging this power-law relationship when using FM can aid in more efficiently allocating resources and optimizing model performance.